# Parabolic induction in the homotopy category of pro-$p$ Iwahori-Hecke modules

### v2

@@ 3.1. Parabolic induction and recollement: one parabolic
Then, in the above setting of a recollement, the category ${\mathcal{T}}^{\prime\prime}$ is equivalent to both $({\mathcal{T}}^{\prime})^{\ddagger}$ and ${}^{\ddagger}({\mathcal{T}}^{\prime})$ via the functors $\mathsf{P}$ and $\mathsf{Q}$ respectively, where we view ${\mathcal{T}}^{\prime}$ as a full subcategory of ${\mathcal{T}}$ via the functor $\mathsf{G}$. Note that by the adjunctions ${\mathsf{L}}\dashv{\mathsf{G}}\dashv{\mathsf{R}}$, we have $({\mathcal{T}}^{\prime})^{\ddagger}={\operatorname{Ker}}({\mathsf{R}})$ and ${}^{\ddagger}({\mathcal{T}}^{\prime})={\operatorname{Ker}}({\mathsf{L}})$.
@@
  1. (i)
    There is an adjoint triple
    ${{\operatorname{Ho}}(S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\operatorname{Ho}}({\mathsf{G}})}$${{\operatorname{Ho}}(R)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\operatorname{Ho}}({\mathsf{L}})}$$\scriptstyle{\mathbf{R}{\mathsf{R}}}$
    and ${\operatorname{Ho}}({\mathsf{G}})$ is fully faithful.
  2. (ii)
    All the above functors are triangulated.
  3. (iii)
    The essential image of ${\operatorname{Ho}}({\mathsf{G}})$ is a localising subcategory of ${\operatorname{Ho}}(R)$.
Proof.
(i) We view our homotopy categories as coming from the Gorenstein injective model structures. Then the existence of such an adjoint triple is immediate from 3.4(ii) and 3.5(i), and 3.5(ii) says that ${\operatorname{Ho}}({\mathsf{G}})$ is fully faithful.
(ii) Recall that for a Gorenstein projective module ${\mathfrak{m}}\in{\operatorname{Mod}}(S)$, its suspension in ${\operatorname{Ho}}(S)$ is defined to be the cokernel of a cofibration to a projective module. Being an exact left Quillen functor with exact right adjoint, the functor ${\mathsf{L}}$ preserves this construction so that ${\operatorname{Ho}}({\mathsf{L}})\circ\Sigma\cong\Sigma\circ{\operatorname{Ho}}({\mathsf{L}})$. Taking successive right adjoints of this identity gives us that all three functors commute with the shift functor. Furthermore, since ${\mathsf{L}}$ is exact and preserves the construction of $\Sigma$, we see that it in fact preserves the distinguished triangles as given in (2.8). Thus ${\operatorname{Ho}}({\mathsf{L}})$ is triangulated. A completely dual argument using fiber sequences and the loop functor shows that ${\operatorname{Ho}}({\mathsf{G}})$ and $\mathbf{R}{\mathsf{R}}$ preserve distinguished triangles, recalling that ${\mathsf{R}}$ preserves short exact sequences of Gorenstein injective modules by 3.7.
(iii) This is immediate from the fact that ${\operatorname{Ho}}({\mathsf{G}})$ commutes with arbitrary direct sums as it is a left adjoint. ∎
Remark 3.11.
Part (ii) of the above proposition also follows immediately from general properties of stable model categories. Indeed, if $({\mathsf{F}},{\mathsf{G}})$ is a Quillen adjunction between stable model categories, then the derived functors $\mathbf{L}{\mathsf{F}}$ and $\mathbf{R}{\mathsf{G}}$ commute with the suspension by [6, Corollary 3.1.4], noting that the loop functor is the inverse of suspension. Moreover, in any such model category the distinguished triangles in the homotopy category are given by cofiber sequences, and left/right derived functors preserve those (c.f. [14, Proposition 6.4.1 & Theorem 7.1.11]).
Corollary 3.12.
The adjoint triple from 3.10(i) is part of a recollement
${{\operatorname{Ho}}(S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\operatorname{Ho}}({\mathsf{G}})}$${{\operatorname{Ho}}(R)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathsf{\pi}}$$\scriptstyle{{\operatorname{Ho}}({\mathsf{L}})}$$\scriptstyle{\mathbf{R}{\mathsf{R}}}$${{\operatorname{Ho}}(R)/{\operatorname{Ho}}(S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathsf{Q}}$$\scriptstyle{\mathsf{P}}$
@@
By 3.10, the inclusion ${\operatorname{Im}}({\operatorname{Ho}}({\mathsf{G}}))\to{\operatorname{Ho}}(R)$ has both a left and a right adjoint and ${\operatorname{Im}}({\operatorname{Ho}}({\mathsf{G}}))$ is a thick subcategory. Then [20, Propositions 4.9.1 & 4.13.1] show that the quotient functor $\pi$ also has a left and a right adjoint and that these two adjoint triples form a recollement as required. ∎
Note that both ${\operatorname{Im}}(\mathsf{Q})={{}^{\ddagger}}{\operatorname{Ho}}(S)$ and ${\operatorname{Ho}}(S)$ are in fact localising subcategories of ${\operatorname{Ho}}(R)$ since they are kernels of triangulated functors which are left adjoints.
@@ 3.2. Extension to several parabolics
with $\ell^{G}_{M}$ and $r^{G}_{M}$ both fully faithful, as given by the recollement in 3.12. The following result will be our main aim in this subsection.
@@
To prove 3.13 we will argue by induction on $n:=\left|{\mathcal{M}}\right|$, the case $n=1$ being 3.12. We will need to establish some preparatory results first, under the assumption that the statement of 3.13(i) holds for some fixed $n\geq 1$ and all split reductive groups. Assuming this from now on, we let =${\mathcal{M}}=\{M_{1},\ldots,M_{n}\}$ be a collection of $n$ distinct standard Levi subgroups and we fix another standard Levi subgroup $L$. Then $L_{i}:=M_{i}\cap L$ is a standard Levi subgroup of $L$, and we may embed fully faithfully each ${\operatorname{Ho}}({\mathcal{H}}_{L_{i}})$ inside ${\operatorname{Ho}}({\mathcal{H}}_{L})$ via ${\operatorname{Ho}}(I_{L_{i}}^{L})$. Thus we may consider the smallest thick subcategory ${\mathcal{S}}_{{\mathcal{M}}\cap L}^{L}$ of ${\operatorname{Ho}}({\mathcal{H}}_{L})$ which contains these essential images, and by assumption we also have a recollement
@@
[…]
@@
(i) By transitivity of parabolic induction (c.f.  [29, Proposition 4.3]), for each $1\leq i\leq n$ we have
@@
Next, we show that the essential image of $(I^{G}_{L})_{\mathcal{M}}$ is localising (and in particular thick). For this, note that since ${\operatorname{Ho}}({\mathcal{H}}_{M})$ and ${\operatorname{Ho}}({\mathcal{H}})$ have arbitrary direct sums so do their quotient categories ${\operatorname{Ho}}({\mathcal{H}}_{L})_{{\mathcal{M}}\cap L}$ and ${\operatorname{Ho}}({\mathcal{H}})_{\mathcal{M}}$ because the quotient functors are left adjoints. Then $(I^{G}_{L})_{\mathcal{M}}$ preserve these direct sums since it is also a left adjoint.
Therefore, we are in the same situation as the one obtained in 3.10, and we may again apply [20, Propositions 4.9.1 & 4.13.1] to obtain a recollement. ∎
@@
We then define the functor $\pi^{G}_{{\mathcal{M}}^{\prime}}$ to be the composite $\pi^{\mathcal{M}}_{{\mathcal{M}}^{\prime}}\circ\pi^{G}_{\mathcal{M}}:{\operatorname{Ho}}({\mathcal{H}})\to{\mathcal{T}}_{{\mathcal{M}}^{\prime}}$ and let ${\mathcal{S}}_{{\mathcal{M}}^{\prime}}:={\operatorname{Ker}}(\pi^{G}_{{\mathcal{M}}^{\prime}})$. Note that $\pi^{G}_{{\mathcal{M}}^{\prime}}$ has fully faithful left and right adjoints, given by $\ell^{G}_{{\mathcal{M}}^{\prime}}:=\ell^{G}_{\mathcal{M}}\circ\ell^{\mathcal{M}}_{{\mathcal{M}}^{\prime}}$ and $r^{G}_{{\mathcal{M}}^{\prime}}:=r^{G}_{\mathcal{M}}\circ r^{\mathcal{M}}_{{\mathcal{M}}^{\prime}}$ respectively. Therefore it follows from [20, Proposition 4.9.1] and its dual that the inclusion ${\mathcal{S}}_{{\mathcal{M}}^{\prime}}\to{\operatorname{Ho}}({\mathcal{H}})$ also has both a left and a right adjoint, and and so these adjunctions form a recollement. Note also that ${\mathcal{S}}_{{\mathcal{M}}^{\prime}}$ is automatically localising, because the functor $\pi^{G}_{{\mathcal{M}}^{\prime}}$ commutes with arbitrary direct sums as it is a left adjoint. We are thus only left to check that ${\mathcal{S}}^{G}_{{\mathcal{M}}^{\prime}}={\mathcal{S}}_{{\mathcal{M}}^{\prime}}$.
@@
For (ii), we assume $n>1$ and fix $1\leq i\leq n$. We denote by $\eta^{G}_{\mathcal{M}}$ (resp. $\eta^{G}_{{\mathcal{M}}_{i}}$, resp. $\eta^{{\mathcal{M}}_{i}}_{\mathcal{M}}$) the unit of the adjunction $\pi^{G}_{{\mathcal{M}}}\dashv r^{G}_{{\mathcal{M}}}$ (resp. $\pi^{G}_{{\mathcal{M}}_{i}}\dashv r^{G}_{{\mathcal{M}}_{i}}$, resp. $\pi^{{\mathcal{M}}_{i}}_{\mathcal{M}}\dashv r^{{\mathcal{M}}_{i}}_{\mathcal{M}}$). Then by the recollement in (i) we have two distinguished triangles
@@
for any ${\mathfrak{m}}\in{\operatorname{Ho}}({\mathcal{H}})$. Recall that $\pi^{G}_{{\mathcal{M}}}=\pi^{{\mathcal{M}}_{i}}_{\mathcal{M}}\circ\pi^{G}_{{\mathcal{M}}_{i}}$ and $r^{G}_{{\mathcal{M}}}=r^{G}_{{\mathcal{M}}_{i}}\circ r^{{\mathcal{M}}_{i}}_{\mathcal{M}}$. By considering the recollement (3.5) for the collection of Levi subgroups ${\mathcal{M}}_{i}$ and $L=M_{i}$ instead, and applying $r^{G}_{{\mathcal{M}}_{i}}$ to the analogous triangle for $\overline{{\mathfrak{m}}}:=\pi^{G}_{{\mathcal{M}}_{i}}({\mathfrak{m}})$, we get
@@
c.f. [5, 0145]. Since $\mathfrak{n}\cong r^{G}_{{\mathcal{M}}_{i}}\pi^{G}_{{\mathcal{M}}_{i}}{\operatorname{Ho}}(I^{G}_{M_{i}})\mathbf{R}\mathcal{R}^{G}_{M_{i}}({\mathfrak{m}})$, c.f. 3.15(i), this is the desired triangle. The construction for the left adjoints is completely dual to this.
Finally, for (iii) we only prove the formula for the essential image of $r^{G}_{\mathcal{M}}$, the proof for $\ell^{G}_{\mathcal{M}}$ being completely analogous. We argue again by induction on $n$, the case $n=1$ being immediate from 3.12. Assume the equality holds for some $n\geq 1$ and any collection of $n$ standard Levi subgroups. Again, let $M_{1},\ldots,M_{n+1}$ be distinct standard Levi subgroups and write =${\mathcal{M}}=\{M_{1},\ldots,M_{n}\}$ and =${\mathcal{M}}^{\prime}={\mathcal{M}}\cup\{M_{n+1}\}$. Note first that ${\operatorname{Im}}(r^{G}_{{\mathcal{M}}^{\prime}})={\operatorname{Ker}}(\mathcal{R}^{G}_{{\mathcal{M}}^{\prime}})$ and ${\operatorname{Im}}(r^{G}_{\mathcal{M}})={\operatorname{Ker}}(\mathcal{R}^{G}_{\mathcal{M}})$. So by the induction hypothesis, one inclusion is automatic from the triangle (3.6) applied to ${\mathcal{M}}^{\prime}$ and $i=n+1$. Assuming now that ${\mathfrak{m}}\in{\operatorname{Ker}}(\mathcal{R}^{G}_{{\mathcal{M}}^{\prime}})$, the same application of (3.6) then implies that $i^{G}_{\mathcal{M}}\mathcal{R}^{G}_{\mathcal{M}}({\mathfrak{m}})$ lies in ${\operatorname{Im}}(r^{G}_{\mathcal{M}})={\operatorname{Ker}}(\mathcal{R}^{G}_{\mathcal{M}})$. But since $i^{G}_{\mathcal{M}}$ is fully faithful we may apply 2.4 to obtain

### v4

@@ 3.1. Parabolic induction and recollement: one parabolic
Then, in the above setting of a recollement, the category ${\mathcal{T}}^{\prime\prime}$ is equivalent to both $({\mathcal{T}}^{\prime})^{\ddagger}$ and ${}^{\ddagger}({\mathcal{T}}^{\prime})$ via the functors $\mathsf{P}$ and $\mathsf{Q}$ respectively, where we view ${\mathcal{T}}^{\prime}$ as a full subcategory of ${\mathcal{T}}$ via the functor $\mathsf{G}$. Note that by the adjunctions ${\mathsf{L}}\dashv{\mathsf{G}}\dashv{\mathsf{R}}$, we have $({\mathcal{T}}^{\prime})^{\ddagger}={\operatorname{Ker}}({\mathsf{R}})$ and ${}^{\ddagger}({\mathcal{T}}^{\prime})={\operatorname{Ker}}({\mathsf{L}})$. In order to establish the existence of a recollement, we have the following:
Proposition 3.10 ([22, Propositions 4.9.1 & 4.13.1]).
Let ${\mathcal{T}}$ be a triangulated category and ${\mathcal{S}}$ a thick subcategory of ${\mathcal{T}}$.
  1. (i)
    The following are equivalent:
    1. (a)
      the inclusion ${\mathcal{S}}\xrightarrow{\subseteq}{\mathcal{T}}$ has a right adjoint;
    2. (b)
      the quotient functor ${\mathcal{T}}\to{\mathcal{T}}/{\mathcal{S}}$ has a a right adjoint; and
    3. (c)
      the inclusion ${\mathcal{S}}^{\ddagger}\xrightarrow{\subseteq}{\mathcal{T}}$ has a left adjoint and ${{}^{\ddagger}}({\mathcal{S}}^{\ddagger})={\mathcal{S}}$.
  2. (ii)
    If the equivalent conditions of (i) and their dual hold, then the inclusion ${\mathcal{S}}\xrightarrow{\subseteq}{\mathcal{T}}$ is part of a recollement
[…]
    of triangulated categories.
We now go back to 3.1. We remind the reader that ${\operatorname{Ho}}(R)$ and ${\operatorname{Ho}}(S)$ as well as their triangulated structure do not depend on which model structure (Gorenstein projective or Gorenstein injective) we work with, cf. 2.12. Applying the above, we now get:
@@
The functors ${\mathsf{G}}$, ${\mathsf{L}}$ and ${\mathsf{R}}$ induce a recollement
${{\operatorname{Ho}}(S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\operatorname{Ho}}({\mathsf{G}})}$${{\operatorname{Ho}}(R)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathsf{\pi}}$$\scriptstyle{{\operatorname{Ho}}({\mathsf{L}})}$$\scriptstyle{\mathbf{R}{\mathsf{R}}}$${{\operatorname{Ho}}(R)/{\operatorname{Ho}}(S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@
We first view our homotopy categories as coming from the Gorenstein injective model structures. Then the existence of the adjoint triple ${\operatorname{Ho}}({\mathsf{L}})\dashv{\operatorname{Ho}}({\mathsf{G}})\dashv\mathbf{R}{\mathsf{R}}$ is immediate from 3.4(ii) and 3.5(i), and 3.5(ii) says that ${\operatorname{Ho}}({\mathsf{G}})$ is fully faithful. We will be done by 3.10(ii) if we show that these functors are triangulated and ${\operatorname{Im}}({\operatorname{Ho}}({\mathsf{G}}))$ is a thick subcategory.
For this, unravelling the definitions, recall that the suspension in ${\operatorname{Ho}}(S)$ of an arbitrary module ${\mathfrak{m}}\in{\operatorname{Mod}}(S)$ (which is automatically cofibrant for the Gorenstein injective model structure) is given as the cokernel of a monomorphism to an injective module. Being an exact functor with exact left adjoint, the functor ${\mathsf{G}}$ preserves this construction so that ${\operatorname{Ho}}({\mathsf{G}})\circ\Sigma\cong\Sigma\circ{\operatorname{Ho}}({\mathsf{G}})$. Taking left and right adjoints of this identity gives us that all three functors commute with the shift functor. Furthermore, since ${\mathsf{G}}$ is exact and preserves the construction of $\Sigma$, we see that it in fact preserves the distinguished triangles as given in (2.11). Thus ${\operatorname{Ho}}({\mathsf{G}})$ is triangulated. A completely dual argument using fiber sequences and the loop functor shows that $\mathbf{R}{\mathsf{R}}$ is also triangulated, recalling that ${\mathsf{R}}$ preserves short exact sequences of Gorenstein injectives by 3.7. To show that ${\operatorname{Ho}}({\mathsf{L}})$ preserves distinguished triangles, we may argue exactly as above but using the Gorenstein projective model structure instead to explicitly describe the triangles. Indeed, ${\mathsf{L}}$ is exact and preserves the explicit construction of the shift, as the cokernel of a cofibration from a Gorenstein projective to a projective module, given by that model structure so that it preserves the construction in (2.11). Finally, ${\operatorname{Im}}({\operatorname{Ho}}({\mathsf{G}}))$ is in fact a localising subcategory and hence is thick, since ${\operatorname{Ho}}({\mathsf{G}})$ commutes with arbitrary direct sums as it is a left adjoint. ∎
Remark 3.12.
The fact that ${\operatorname{Ho}}({\mathsf{L}})$, ${\operatorname{Ho}}({\mathsf{G}})$ and $\mathbf{R}{\mathsf{R}}$ are triangulated also follows immediately from general properties of stable model categories. Indeed, if $({\mathsf{F}},{\mathsf{G}})$ is a Quillen adjunction between stable model categories, then the derived functors $\mathbf{L}{\mathsf{F}}$ and $\mathbf{R}{\mathsf{G}}$ commute with the suspension by [6, Corollary 3.1.4], noting that the loop functor is the inverse of suspension. Moreover, in any such model category the distinguished triangles in the homotopy category are given by cofiber sequences, and left/right derived functors preserve those (cf. [16, Proposition 6.4.1 & Theorem 7.1.11]).
@@ 3.2. Extension to several parabolics
with $\ell^{G}_{M}$ and $r^{G}_{M}$ both fully faithful, as given by the recollement in 3.11. The following result will be our main aim in this subsection.
@@
To prove 3.13 we will argue by induction on $n:=\left|{\mathcal{M}}\right|$, the case $n=1$ being 3.11. We will need to establish some preparatory results first, under the assumption that the statement of 3.13(i) holds for some fixed $n\geq 1$ and all split reductive groups. Assuming this from now on, we let =${\mathcal{M}}=\{M_{1},\ldots,M_{n}\}$ be a collection of $n$ distinct standard Levi subgroups and we fix another standard Levi subgroup $L$. Then $L_{i}:=M_{i}\cap L$ is a standard Levi subgroup of $L$, and we may embed fully faithfully each ${\operatorname{Ho}}({\mathcal{H}}_{L_{i}})$ inside ${\operatorname{Ho}}({\mathcal{H}}_{L})$ via ${\operatorname{Ho}}(I_{L_{i}}^{L})$. Thus we may consider the smallest thick subcategory ${\mathcal{S}}_{{\mathcal{M}}\cap L}^{L}$ of ${\operatorname{Ho}}({\mathcal{H}}_{L})$ which contains these essential images, and by assumption we also have a recollement
@@
    $\text{and}\quad\quad\pi_{{\mathcal{M}}\cap L}^{L}\circ\mathbf{R}\mathcal{R}_{L}^{G}$ $=(\mathcal{R}^{G}_{L})_{\mathcal{M}}\circ\pi_{\mathcal{M}}^{G}.$
@@
(i) By transitivity of parabolic induction (cf.  [32, Proposition 4.3]), for each $1\leq i\leq n$ we have
@@
Finally, we will be done by 3.10(ii) if we show that the essential image of $(I^{G}_{L})_{\mathcal{M}}$ is localising (and in particular thick). For this, note that since ${\operatorname{Ho}}({\mathcal{H}}_{M})$ and ${\operatorname{Ho}}({\mathcal{H}})$ have arbitrary direct sums so do their quotient categories ${\operatorname{Ho}}({\mathcal{H}}_{L})_{{\mathcal{M}}\cap L}$ and ${\operatorname{Ho}}({\mathcal{H}})_{\mathcal{M}}$ because the quotient functors are left adjoints. Then $(I^{G}_{L})_{\mathcal{M}}$ preserves these direct sums since it is also a left adjoint. ∎
@@
We then define the functor $\pi^{G}_{{\mathcal{M}}^{\prime}}$ to be the composite $\pi^{\mathcal{M}}_{{\mathcal{M}}^{\prime}}\circ\pi^{G}_{\mathcal{M}}:{\operatorname{Ho}}({\mathcal{H}})\to{\mathcal{T}}_{{\mathcal{M}}^{\prime}}$ and let ${\mathcal{S}}_{{\mathcal{M}}^{\prime}}:={\operatorname{Ker}}(\pi^{G}_{{\mathcal{M}}^{\prime}})$. Note that $\pi^{G}_{{\mathcal{M}}^{\prime}}$ has fully faithful left and right adjoints, given by $\ell^{G}_{{\mathcal{M}}^{\prime}}:=\ell^{G}_{\mathcal{M}}\circ\ell^{\mathcal{M}}_{{\mathcal{M}}^{\prime}}$ and $r^{G}_{{\mathcal{M}}^{\prime}}:=r^{G}_{\mathcal{M}}\circ r^{\mathcal{M}}_{{\mathcal{M}}^{\prime}}$ respectively. Therefore it follows from 3.10 that the inclusion ${\mathcal{S}}_{{\mathcal{M}}^{\prime}}\to{\operatorname{Ho}}({\mathcal{H}})$ also has both a left and a right adjoint, and that these adjunctions form a recollement. Note also that ${\mathcal{S}}_{{\mathcal{M}}^{\prime}}$ is automatically localising, because the functor $\pi^{G}_{{\mathcal{M}}^{\prime}}$ commutes with arbitrary direct sums as it is a left adjoint. We are thus only left to check that ${\mathcal{S}}^{G}_{{\mathcal{M}}^{\prime}}={\mathcal{S}}_{{\mathcal{M}}^{\prime}}$.
@@
For (ii), we assume $n>1$ and fix $1\leq i\leq n$. We denote by $\eta^{G}_{\mathcal{M}}$ (resp. $\eta^{G}_{{\mathcal{M}}_{i}}$, resp. $\eta^{{\mathcal{M}}_{i}}_{\mathcal{M}}$) the unit of the adjunction $\pi^{G}_{{\mathcal{M}}}\dashv r^{G}_{{\mathcal{M}}}$ (resp. $\pi^{G}_{{\mathcal{M}}_{i}}\dashv r^{G}_{{\mathcal{M}}_{i}}$, resp. $\pi^{{\mathcal{M}}_{i}}_{\mathcal{M}}\dashv r^{{\mathcal{M}}_{i}}_{\mathcal{M}}$). Then by the recollement in (i) and (3.5) we have two distinguished triangles
@@
for any ${\mathfrak{m}}\in{\operatorname{Ho}}({\mathcal{H}})$. Recall that $\pi^{G}_{{\mathcal{M}}}=\pi^{{\mathcal{M}}_{i}}_{\mathcal{M}}\circ\pi^{G}_{{\mathcal{M}}_{i}}$ and $r^{G}_{{\mathcal{M}}}=r^{G}_{{\mathcal{M}}_{i}}\circ r^{{\mathcal{M}}_{i}}_{\mathcal{M}}$. Next, we consider the triangle (3.5) for the object $\overline{{\mathfrak{m}}}:=\pi^{G}_{{\mathcal{M}}_{i}}({\mathfrak{m}})$ in the setting of the recollement (3.6) for the collection of Levi subgroups ${\mathcal{M}}_{i}$ and $L=M_{i}$. Applying the triangulated functor $r^{G}_{{\mathcal{M}}_{i}}$ to it, we get
@@
cf. [5, 0145]. Since $\mathfrak{n}\cong r^{G}_{{\mathcal{M}}_{i}}\pi^{G}_{{\mathcal{M}}_{i}}{\operatorname{Ho}}(I^{G}_{M_{i}})\mathbf{R}\mathcal{R}^{G}_{M_{i}}({\mathfrak{m}})$, cf. 3.15(i), this is the desired triangle. The construction for the left adjoints is completely dual to this.
Finally, for (iii) we only prove the formula for the essential image of $r^{G}_{\mathcal{M}}$, the proof for $\ell^{G}_{\mathcal{M}}$ being completely analogous. We argue again by induction on $n$, the case $n=1$ being immediate from 3.11. Assume the equality holds for some $n\geq 1$ and any collection of $n$ standard Levi subgroups. Again, let $M_{1},\ldots,M_{n+1}$ be distinct standard Levi subgroups and write =${\mathcal{M}}=\{M_{1},\ldots,M_{n}\}$ and =${\mathcal{M}}^{\prime}={\mathcal{M}}\cup\{M_{n+1}\}$. Note first that ${\operatorname{Im}}(r^{G}_{{\mathcal{M}}^{\prime}})={\operatorname{Ker}}(\mathcal{R}^{G}_{{\mathcal{M}}^{\prime}})$ and ${\operatorname{Im}}(r^{G}_{\mathcal{M}})={\operatorname{Ker}}(\mathcal{R}^{G}_{\mathcal{M}})$. So by the induction hypothesis, the inclusion $\bigcap_{i=1}^{n+1}{\operatorname{Ker}}(\mathbf{R}\mathcal{R}^{G}_{M_{i}})\subseteq{\operatorname{Im}}(r^{G}_{{\mathcal{M}}^{\prime}})$ is automatic from the triangle (3.7) applied to ${\mathcal{M}}^{\prime}$ and $i=n+1$. Assuming now that ${\mathfrak{m}}\in{\operatorname{Ker}}(\mathcal{R}^{G}_{{\mathcal{M}}^{\prime}})$, the same application of (3.7) then implies that $i^{G}_{\mathcal{M}}\mathcal{R}^{G}_{\mathcal{M}}({\mathfrak{m}})$ lies in ${\operatorname{Im}}(r^{G}_{\mathcal{M}})={\operatorname{Ker}}(\mathcal{R}^{G}_{\mathcal{M}})$. But since $i^{G}_{\mathcal{M}}$ is fully faithful we may apply 2.4 to obtain